What is the surface area of the square pyramid with an 11 inches square base and a 15 inches slant height of the pyramid?
Answer by checkley71(8403)

HERE YOU HAVE 4 TRIANGLES PLUS A SQUARE BASE.
THE SQUARE BASE=11*11=121 IN^2
THE 4 TRIANGLE SIDES=4*11/2*15=330 IN^2
NOW WE ADD THESE 2 FIGURES
121+330=451 IN^2